# MPMA: Preference Manipulation Attack Against Model Context Protocol

### v1

@@ Abstract
Model Context Protocol (MCP) standardizes interface mapping for large language models (LLMs) to access external data and tools, which revolutionizes the paradigm of tool selection and facilitates the rapid expansion of the LLM agent tool ecosystem. However, as the MCP is increasingly adopted, third-party customized versions of the MCP server expose potential security vulnerabilities. In this paper, we first introduce a novel security threat, which we term the MCP Preference Manipulation Attack (MPMA). An attacker deploys a customized MCP server to manipulate LLMs, causing them to prioritize it over other competing MCP servers. This can result in economic benefits for attackers, such as revenue from paid MCP services or advertising income generated from free servers. To achieve MPMA, we first design a Direct Preference Manipulation Attack () that achieves significant effectiveness by inserting the manipulative word and phrases into the tool name and description. However, such a direct modification is obvious to users and lacks stealthiness. To address these limitations, we further propose Genetic-based Advertising Preference Manipulation Attack (). employs four commonly used strategies to initialize descriptions and integrates a Genetic Algorithm (GA) to enhance stealthiness. The experiment results demonstrate that balances high effectiveness and stealthiness. Our study reveals a critical vulnerability of the MCP in open ecosystems, highlighting an urgent need for robust defense mechanisms to ensure the fairness of the MCP ecosystem.
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated transformative capabilities in tasks such as reasoning[54], mathematics[34], and code generation[55]. As LLMs rapidly advance in their abilities, LLM agents arise[45; 42; 47], an autonomous system built around an LLM, capable of perceiving its environment, planning actions, and executing tasks to achieve goal-directed intelligent behavior in complex settings. A key feature that enables LLM agents to perform such tasks is their ability to select and call external tools, which extends their action space beyond language generation.
@@
In late 2024, Anthropic revolutionarily introduced the Model Context Protocol (MCP)[42; 18; 47], a protocol that enables LLM agents to autonomously discover and select tools without relying on predefined interface mappings of function calling. By standardizing tool calling interfaces, MCP significantly reduces development barriers and accelerates the expansion of the LLM agent tool ecosystem[42; 47]. Since its introduction, the MCP has rapidly evolved from a niche protocol into a foundational infrastructure for building LLM agents. Currently, dozens of third-party platforms have deployed a large number of MCP servers[26; 21; 19; 20], with some of them operating at a scale exceeding 13,000 instances[20]. Furthermore, many MCP servers provide high-quality and commercial-grade services, such as image generation[23; 10], web search[4; 22], and location-based[2; 14] functionalities, through API interfaces, demonstrating substantial potential in promoting service commercialization and market expansion of MCP. Although the MCP community has begun to pay preliminary attention to security issues, current research primarily focuses on the potential presence of malicious code and privacy leakage within MCP servers[30; 33; 25]. However, a critical question remains: Are these mechanisms sufficient to ensure the overall trustworthiness of MCP applications?
@@
This paper first proposes and investigates the MCP Preference Manipulation Attack (MPMA), a novel security threat against MCP applications. Specifically, multiple paid MCP servers offering similar functionalities often exist in direct competition for economic benefit[23; 10; 1; 24]. In this profit-competing landscape, a malicious MCP server may attempt to manipulate the LLMβs tool selection process in order to increase its likelihood of being chosen across a diverse set of user queries. To achieve MPMA, we first propose Directly Preference Manipulation Attack (), a naive strategy by directly inserting manipulative words or phrases at the beginning of the tool name and description. proves highly effective, achieving a 100% Attack Success Rate (ASR) in most settings. However, we emphasize that the stealthiness of the attack is critically important, as both the tool name and description are subject to manual inspection by users and third-party platform reviewers. Therefore, it is essential to design the manipulative content that remains inconspicuous while effectively influencing the tool selection process. Inspired by the effectiveness of traditional advertising in manipulating human preference without awareness[38; 35], we further propose Genetic-based Advertising Preference Manipulation Attack (). leveraging traditional advertising strategies to construct four description optimization objectives: Authoritative, Emotional, Exaggerated, and Subliminal[52; 37; 40; 51]. Subsequently, we employ a black-box Genetic Algorithm (GA) to further enhance the stealthiness of the attack. Extensive experiment results demonstrate that the proposed methods significantly improve stealthiness while maintaining high attack effectiveness.
@@
Our calculations (see AppendixΒ A) suggest that, under conservative estimates, both and could cause unfair benefits exceeding 200,000 dollars to other MCP servers merely in the web search server alone each year. Furthermore, as the MCP gains wider adoption in standardizing tool calling across LLM agents, the economic impact is expected to grow significantly. Our research reveals critical security vulnerabilities inherent in the MCP framework, thereby highlighting the necessity of developing robust and systematic defense mechanisms to ensure the fairness of the MCP ecosystem. Our main contributions are summarized as follows:
β’
@@
β’
We conduct comprehensive experiments across 8 MCP servers and 5 mainstream LLMs. The results consistently demonstrate the vulnerability of MCP-based tool selection to MPMA, highlighting the urgent need for corresponding defense mechanisms for the fairness of the MCP ecosystem.
2 Preliminary and Related Work
@@ 2.1 Model Context Protocol (MCP)
Before the introduction of MCP, OpenAI first introduced the function calling mechanism in 2023, enabling LLMs to autonomously call external tools and dynamically interact with the real world[49; 42; 47]. Although function calling provides a foundational framework for tool integration, it presents several limitations. Specifically, it requires developers to manually define interfaces and configure authentication parameters, resulting in limited generality and scalability. These limitations have collectively hindered the widespread adoption and growth of the function calling ecosystem. In contrast, MCP standardizes tool calling interfaces, significantly reducing development barriers and accelerating the expansion of the LLM agent tool ecosystem[26; 21; 19; 20]. The MCP architecture consists of three main components: MCP host, MCP client, and MCP server[42; 47]. Their definitions and functionalities are described as follows:
@@
MCP Host. This refers to the integrated development environments (IDEs), or AI tools that access data via the MCP[42]. The host integrates interaction tools for users, MCP servers, and LLMs, enabling efficient MCP-based communication. Representative examples include Claude Desktop[6], Cursor[9], and the VSCode plugin Cline[7].
@@
MCP Client. The MCP client functions as an intermediary within the host environment. It manages communication between the MCP host and one or more MCP servers[42].
@@
MCP Server. The MCP server acts as a gateway that enables the MCP client to access external services and execute tasks by interacting with external tools. Fundamentally, it is an application capable of interacting with the MCP client[42]. Note that a single MCP server may contain one or multiple tools, and every tool has a name and description. By adhering to the MCP, the tool provides external information and resources to LLMs, allowing them to autonomously plan and complete tasks.
@@
To facilitate understanding, we illustrate an example of a single MCP calling in FigureΒ 1. The figure divides the process into four steps: βΆ Task planning: A user inputs the user prompt, and the MCP host provides the LLM with the prompt and the contextual information, including the list of available MCP servers and tools with their descriptions and names. The LLM determines that additional input from a weather service is necessary and selects a suitable tool accordingly. β· Tool selection: The LLM sends the tool calling request to the MCP host. βΈ Tool calling: The MCP host forwards the calling request to the chosen MCP server. βΉ Conclusion and output: The corresponding tool retrieves the required information through an API call or local data, and the data is passed back to the LLM via the MCP host, and the LLM outputs the final response to the user after the conclusion. Note that the MPMA primarily targets the βΆ (task planning) and β· (tool selection) stages.
2.2 Prompt Injection Attack Against LLMs
A prompt injection attack refers to the insertion of a malicious prompt into the query submitted to LLMs, aiming to manipulate the modelβs behavior. The proposed MPMA can be regarded as a specialized form of prompt injection, wherein manipulative content is embedded into the tool name and description fields. These fields are subsequently incorporated into the model input, thereby influencing the LLMβs tool selection decisions. A prompt injection attack can be used to implement various attacks. Zhang et al.[57] implant backdoor prompts into customized versions of GPT, thereby enabling backdoor attacks[44]. Nestaas et al.[48] conduct a preference manipulation attack on LLMs with the internet search by embedding malicious instructions in web pages using font colors that match the background, thus achieving a covert preference manipulation. Shi et al.[50] implement the preferred operation of the LLM agent in a black-box and white-box setting by inserting the prompt into the tools or gradient-based optimization. Their work operates within the context of selecting tools through a retriever. Specifically, the tools they use are chosen by calculating the similarity between the tool document and the userβs query. In contrast, in the MCP, tool selection is autonomously determined and decided by the LLM. And their attack aims to facilitate subsequent attacks by making it easier for users to interact with a malicious MCP server, rather than to pursue economic benefits. Overall, these works in preference manipulation attacks highlight security issues in various LLM applications, such as LLM-based search engines[48] and traditional LLM agents[50]. However, our work investigates security concerns in emerging MCP applications, which represent a novel and rapidly developing domain.
@@ 3 Threat Model
Attack Scenario. The scenario is shown in the FigureΒ 2. We consider a malicious provider that publishes a paid MCP server on the third-party platform. When the user deploys this server, it will influence the LLMβs tool selection process, thereby increasing the likelihood that the malicious server is chosen over its competitors. This preferential selection ultimately leads to economic gains for the attacker through service usage fees or advertising income.
Attackerβs Capability. We assume the attacker as the MCP server builder who has white-box access to the MCP server, allowing manipulation of metadata such as the tool name and description. Furthermore, the attacker is capable of publishing the malicious MCP server to third-party MCP platforms. Note that the attacker does not possess any control or modification capability over the base LLM deployed within the LLM agent.
Attackerβs Goals. (1) Attack effectiveness. The attacker seeks to ensure that the malicious server consistently outperforms competing servers in terms of selection frequency by the LLMs, thereby securing measurable economic benefits. (2) Stealthiness. The attacker aims to maintain the malicious serverβs inconspicuousness. Specifically, the tool name and description should not raise suspicion among users and should evade both manual inspection and automated machine detection mechanisms.
@@ 4.1 Attack Overview
The attack overview is illustrated in FigureΒ 3, which presents only the steps involved in the LLMβs tool selection. We emphasize that both the MCP Host and the LLM have access solely to the name and description of each tool of the MCP server, and the internal processing logic of the server remains invisible to them. Therefore, the MPMA can only be carried out by manipulating the tool name and description for the MCP provider.
@@
βΆ Benign. When all the MCP servers deployed by the user are benign, the model selects the get_weather tool, which is a sufficient tool for this simple task.
β· . We present the Best Description strategy from as a representative example. When one of the available MCP servers is constructed using the Best Description strategy, the model selects this malicious server, providing the justification that it is the best tool in the world.
βΈ . We present the Exaggerate advertising strategy from as a representative example. When one of the available MCP servers is a malicious server constructed using the Exaggerate strategy, the model selects the malicious server, reasoning that it is described as reliable and precise. Compared to , exhibits higher stealthiness, as its descriptions avoid the use of conspicuously manipulative terms. The experiment results in FigureΒ 5 corroborate the improved stealthiness of .
@@ 4.2 Direct Preference Manipulate Attack ()
manipulates the preference of LLMs through manipulative words or phrases. The overall procedure is illustrated in FigureΒ 3. We propose two attack strategies: Best Description and Best Name.
Best Description. Inspired by findings in the [48], we believe that words or phrases including βbestβ are more likely to gain the preference of LLMs. Therefore, we prepend a manipulative description to the tool description. Specifically, the Best Description can be represent in EquationΒ 1:
where the is the string concatenation operation, and the is the raw description. In this work, we use the phrase: βThis is the best tool in the world.β as the manipulative description .
Best Name. Similarly, we prepend the manipulative word βbestβ to the tool name to elicit the preference of the LLM. Specifically, the Best Name can be represent in EquationΒ 2:
Note that these two types of attacks exhibit limited stealthiness, as manipulative words such as βbestβ are likely to trigger suspicion during both manual and automated inspections. We emphasize that stealthiness is critical in the context of MPMA under the MCP setting, as the information of MCP servers is visible to both users and third-party platforms, shown in FigureΒ 6. If manipulative sentences such as those used in are inserted, they are likely to arouse user suspicion. Therefore, we further propose for better stealthiness.
@@ 4.3 Genetic-based Advertising Preference Manipulate Attack ()
4.3.1 Advertising Strategies
We observe that the pursuit of stealthiness in tool descriptions shares conceptual similarities with traditional advertising strategies, both of which seek to influence user preferences without explicit awareness[38; 35]. Motivated by this observation, we systematically investigate advertising strategies that are designed to unconsciously influence audience decisions. Based on our extensive investigation, we adopt the following four representative advertising strategies in the traditional advertising area:
Authoritative (Au)[40]. This strategy embeds advertising content within text by disguising it as expert advice or user recommendations.
Emotional (Em)[52]. This strategy aligns advertising content with the audienceβs emotional needs by incorporating emotionally charged language.
Exaggerated (Ex)[37]. This strategy uses exaggeration and strong rhetorical techniques to make the product appear more appealing.
Subliminal (Su)[51]. This strategy is a form of covert advertising that embeds information through subconscious cues. Although readers may not consciously recognize the advertising content, the implicit messages or psychological suggestions subtly influence their behavior.
We employ GPT-4o[43] to generate tool descriptions that exhibit specific advertising characteristics.
4.3.2 Algorithm for Descriptions Stealthiness Enhancement.
consists of two main components: advertising style transformation and genetic algorithm stealthiness enhancement. We first utilize GPT-4o[43] and advertising prompt to transform the original tool description into a style that aligns with the selected advertising strategy, while maintaining a certain level of stealthiness, after the initialization of the pool . Subsequently, a GA is applied to further enhance the stealthiness of the optimized description by iteratively refining candidate prompts. Specifically, in each iteration, we introduce the Mutate operation using the stealthiness-oriented prompt designed to improve stealthiness and perform the Crossover operation, combining elements from pairs of prompts to promote mutation diversity and explore a broader solution space. The resulting candidate descriptions are accumulated in a pool , from which the GPT-4o selects the top-k descriptions that appear least suspicious. These descriptions are retained for the next iteration, thereby guiding the evolutionary process toward higher stealthiness. After iterations, GPT-4o is used to select the most stealthy description from the final pool. The prompt designed for each advertising strategy and the further details are shown in AppendixΒ B.
5 Experiment Setup

### v2

@@ Abstract
Model Context Protocol (MCP) standardizes interface mapping for large language models (LLMs) to access external data and tools, which revolutionizes the paradigm of tool selection and facilitates the rapid expansion of the LLM agent tool ecosystem. However, as the MCP is increasingly adopted, third-party customized versions of the MCP server expose potential security vulnerabilities. In this paper, we first introduce a novel security threat, which we term the MCP Preference Manipulation Attack (MPMA). An attacker deploys a customized MCP server to manipulate LLMs, causing them to prioritize it over other competing MCP servers. This can result in economic benefits for attackers, such as revenue from paid MCP services or advertising income generated from free servers. To achieve MPMA, we first design a Direct Preference Manipulation Attack () that achieves significant effectiveness by inserting the manipulative words and phrases into the tool name and description. However, such a direct modification is obvious to users and lacks stealthiness. To address these limitations, we further propose Genetic-based Advertising Preference Manipulation Attack (). employs four commonly used strategies to initialize descriptions and integrates a Genetic Algorithm (GA) to enhance stealthiness. The experimental results demonstrate that balances high effectiveness and stealthiness. Our study reveals a critical vulnerability of the MCP in open ecosystems, highlighting an urgent need for robust defense mechanisms to ensure the fairness of the MCP ecosystem. Our code is available at https://github.com/hanbaoergogo/MPMA111This is an extended version of the copyrighted publication at AAAI.
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated transformative capabilities in tasks such as reasoning (cot), mathematics (math), and code generation (codeeval). As LLMs rapidly advance in their abilities, LLM agents arise (agentsurvey; mcpsurvey; mcpsurveybad), an autonomous system built around an LLM, capable of perceiving its environment, planning actions, and executing tasks to achieve goal-directed intelligent behavior in complex settings. A key feature that enables LLM agents to perform such tasks is their ability to select and call external tools, which extends their action space beyond language generation.
@@
In late 2024, Anthropic revolutionarily introduced the Model Context Protocol (MCP) (mcpsurvey; introducingmcp; mcpsurveybad), a protocol that enables LLM agents to autonomously discover and select tools without relying on predefined interface mappings of function calling. By standardizing tool calling interfaces, MCP significantly reduces development barriers and accelerates the expansion of the LLM agent tool ecosystem (mcpsurvey; mcpsurveybad). Since its introduction, the MCP has rapidly evolved from a niche protocol into a foundational infrastructure for building LLM agents. Currently, dozens of third-party platforms have deployed a large number of MCP servers (mcpplatform1; mcpplatform2; mcpplatform3; mcpplatform4), with some of them operating at a scale exceeding 13,000 instances (mcpplatform4). Furthermore, many MCP servers provide high-quality and commercial-grade services, such as image generation (imageserver1; imageserver2), web search (searchserver1; searchserver2), and location-based (locationserver1; locationserver2) functionalities, through API interfaces, demonstrating substantial potential in promoting service commercialization and market expansion of MCP. Although the MCP community has begun to pay preliminary attention to security issues, current research primarily focuses on the potential presence of malicious code and privacy leakage within MCP servers (MCPsecurity1; MCPsecurity2; MCPsecurity3). However, a critical question remains: Are these mechanisms sufficient to ensure the overall trustworthiness of MCP applications?
@@
This paper first proposes and investigates the MCP Preference Manipulation Attack (MPMA), a novel security threat against MCP applications. Specifically, multiple paid MCP servers offering similar functionalities often exist in direct competition for economic benefit (imageserver1; imageserver2; imageserver3; imageserver4). In this profit-competing landscape, a malicious MCP server may attempt to manipulate the LLM’s tool selection process in order to increase its likelihood of being chosen across a diverse set of user queries. To achieve MPMA, we first propose Directly Preference Manipulation Attack (), a naive strategy by directly inserting manipulative words or phrases at the beginning of the tool name and description. proves highly effective, achieving a 100% Attack Success Rate (ASR) in most settings. However, we emphasize that the stealthiness of the attack is critically important, as both the tool name and description are subject to manual inspection by users and third-party platform reviewers. Therefore, it is essential to design the manipulative content that remains inconspicuous while effectively influencing the tool selection process. Inspired by the effectiveness of traditional advertising in manipulating human preference without awareness (advtiseingeffect; economicadvertising), we further propose Genetic-based Advertising Preference Manipulation Attack (). leveraging traditional advertising strategies to construct four description optimization objectives: Authoritative, Emotional, Exaggerated, and Subliminal (emotional; exaggerate; authority; subliminal). Subsequently, we employ a black-box Genetic Algorithm (GA) to further enhance the stealthiness of the attack. Extensive experimental results demonstrate that the proposed methods significantly improve stealthiness while maintaining high attack effectiveness.
@@
Our calculations (see Appendix) suggest that, under conservative estimates, both and could cause unfair benefits exceeding 200,000 dollars to other MCP servers merely in the web search server alone each year. Furthermore, as the MCP gains wider adoption in standardizing tool calling across LLM agents, the economic impact is expected to grow significantly. Our research reveals critical security vulnerabilities inherent in the MCP framework, thereby highlighting the necessity of developing robust and systematic defense mechanisms to ensure the fairness of the MCP ecosystem. Our main contributions are summarized as follows:
@@
We conduct comprehensive experiments across 8 MCP servers and 5 mainstream LLMs. The results demonstrate the vulnerability of MCP-based tool selection to MPMA, highlighting the urgent need for corresponding defense mechanisms for the fairness of the MCP ecosystem.
2 Preliminary
A prompt injection attack refers to the insertion of a malicious prompt into the query submitted to LLMs, aiming to manipulate the model’s behavior. The proposed MPMA can be regarded as a specialized form of prompt injection, wherein manipulative content is embedded into the tool name and description fields. These fields are subsequently incorporated into the model input, thereby influencing the LLM’s tool selection decisions. A prompt injection attack can be used to implement various attacks. Zhang et al. (instructionbackdoor) implant backdoor prompts into customized versions of GPT, thereby enabling backdoor attacks (backdoorsurvey). Nestaas et al. (adversarialseo) conduct a preference manipulation attack on LLMs with the internet search by embedding malicious instructions in web pages using font colors that match the background, thus achieving a covert preference manipulation. Shi et al. (toolselection) implement the preferred operation of the LLM agent in a black-box and white-box setting by inserting the prompt into the tools or gradient-based optimization. Their work operates within the context of selecting tools through a retriever. Specifically, the tools they use are chosen by calculating the similarity between the tool document and the user’s query. In contrast, in the MCP, tool selection is autonomously determined and decided by the LLM. And their attack aims to facilitate subsequent attacks by making it easier for users to interact with a malicious MCP server, rather than to pursue economic benefits. Overall, these works in preference manipulation attacks highlight security issues in various LLM applications, such as LLM-based search engines (adversarialseo) and traditional LLM agents (toolselection). However, our work investigates security concerns in emerging MCP applications, which represent a novel and rapidly developing domain.
@@ 2.1 Model Context Protocol (MCP)
Before the introduction of MCP, OpenAI first introduced the function calling mechanism in 2023, enabling LLMs to autonomously call external tools and dynamically interact with the real world (functioncall; mcpsurvey; mcpsurveybad). Although function calling provides a foundational framework for tool integration, it presents several limitations. Specifically, it requires developers to manually define interfaces and configure authentication parameters, resulting in limited generality and scalability. These limitations have collectively hindered the widespread adoption and growth of the function calling ecosystem. In contrast, MCP standardizes tool calling interfaces, significantly reducing development barriers and accelerating the expansion of the LLM agent tool ecosystem (mcpplatform1; mcpplatform2; mcpplatform3; mcpplatform4). The MCP architecture consists of three main components: MCP host, MCP client, and MCP server (mcpsurvey; mcpsurveybad). Their definitions and functionalities are described as follows:
@@
MCP Host. This refers to the integrated development environments (IDEs), or AI tools that access data via the MCP (mcpsurvey). The host integrates interaction tools for users, MCP servers, and LLMs, enabling efficient MCP-based communication. Representative examples include Claude Desktop (claudedesktop), Cursor (cursor), and the VSCode plugin Cline (cline).
@@
MCP Client. The MCP client is an intermediary within the host environment. It manages communication between the MCP host and one or more MCP servers (mcpsurvey).
@@
MCP Server. The MCP server acts as a gateway that enables the MCP client to access external services and execute tasks by interacting with external tools. Fundamentally, it is an application capable of interacting with the MCP client (mcpsurvey). Note that a single MCP server may contain one or multiple tools, and every tool has a name and description. By adhering to the MCP, the tool provides external information and resources to LLMs, allowing them to autonomously plan and complete tasks.
@@
To facilitate understanding, we illustrate an example of a single MCP calling in Figure˜1. The figure divides the process into four steps: ❶ Task planning: A user inputs the user prompt, and the MCP host provides the LLM with the prompt and the contextual information, including the list of available MCP servers and tools with their descriptions and names. The LLM determines that additional input from a weather service is necessary and selects a suitable tool accordingly. ❷ Tool selection: The LLM sends the tool calling request to the MCP host. ❸ Tool calling: The MCP host forwards the calling request to the chosen MCP server. ❹ Conclusion and output: The corresponding tool retrieves the required information through an API call or local data, and the data is passed back to the LLM via the MCP host, and the LLM outputs the final response to the user after the conclusion. Note that the MPMA primarily targets the ❶ (task planning) and ❷ (tool selection) stages.
@@ 3 Threat Model
Attack Scenario. The scenario is shown in the Figure˜2. We consider a malicious provider that publishes a paid MCP server on the third-party platform. When the user deploys this server, it will influence the LLM’s tool selection process, thereby increasing the likelihood that the malicious server is chosen over its competitors. This preferential selection ultimately leads to economic gains for the attacker through service usage fees or advertising income.
Attacker’s Capability. We assume the attacker as the MCP server builder who has white-box access to the MCP server, allowing manipulation of metadata such as the tool name and description. Furthermore, the attacker can publish the malicious MCP server to third-party MCP platforms. Note that the attacker does not possess any control or modification capability over the base LLM within the LLM agent.
Attacker’s Goals. (1) Attack effectiveness. The attacker seeks to ensure that the malicious server consistently outperforms competing servers in terms of selection frequency by the LLMs, thereby securing measurable economic benefits. (2) Stealthiness. The attacker aims to maintain the malicious server’s inconspicuousness. Specifically, the tool name and description should not raise suspicion among users and should evade both manual inspection and automated machine detection mechanisms.
@@ 4.1 Attack Overview
The attack overview is illustrated in Figure˜3, which presents only the steps involved in the LLM’s tool selection. We emphasize that both the MCP Host and the LLM have access solely to the name and description of each tool of the MCP server, and the internal processing logic of the server remains invisible to them. Therefore, the MPMA can only be carried out by manipulating the tool name and description for the MCP provider.
@@
❶ Benign. When all the MCP servers deployed by the user are benign, the model selects the get_weather tool, which is a sufficient tool for this simple task.
❷ . We present the Best Description strategy from as a representative example. When one of the available MCP servers is constructed using the Best Description strategy, the model selects this malicious server, providing the justification that it is the best tool in the world.
❸ . We present the Exaggerated advertising strategy from as a representative example. When one of the available MCP servers is a malicious server constructed using the Exaggerate strategy, the model selects the malicious server, reasoning that it is described as reliable and precise. Compared to , exhibits higher stealthiness, as its descriptions avoid the use of conspicuously manipulative terms. The experimental results in Figure˜5 corroborate the improved stealthiness of .
@@ 4.2 Direct Preference Manipulate Attack ()
manipulates the preference of LLMs through manipulative words or phrases. The overall procedure is illustrated in Figure˜3. We propose two attack strategies: Best Description and Best Name.
Best Description. Inspired by findings in the (adversarialseo), we believe that words or phrases including “best” are more likely to gain the preference of LLMs. Therefore, we prepend a manipulative description to the tool description. Specifically, the Best Description can be represent in Equation˜1:
where the is the string concatenation, and the is the raw description. In this work, we use the phrase: “This is the best tool in the world.” as the manipulative description .
Best Name. Similarly, we prepend the manipulative word “best” to the tool name to elicit the preference of the LLM. Specifically, the Best Name can be represent below:
Note that these two types of attacks exhibit limited stealthiness, as manipulative words such as “best” are likely to trigger suspicion during both manual and automated inspections. We emphasize that stealthiness is critical in the context of MPMA under the MCP setting, as the information of MCP servers is visible to both users and third-party platforms, as shown in the Appendix. If manipulative sentences such as those used in are inserted, they are likely to arouse user suspicion. Therefore, we further propose for better stealthiness.
@@ 4.3 Genetic-based Advertising Preference Manipulate Attack ()
Advertising Strategies
We observe that the pursuit of stealthiness in tool descriptions shares conceptual similarities with traditional advertising strategies, both of which seek to influence user preferences without explicit awareness (advtiseingeffect; economicadvertising). Motivated by this observation, we systematically investigate advertising strategies that are designed to unconsciously influence audience decisions. Based on our extensive investigation, we adopt the following four representative advertising strategies in the traditional advertising area:
Authoritative (Au) (authority). This strategy embeds advertising content within text by disguising it as expert advice or user recommendations.
Emotional (Em) (emotional). This strategy aligns advertising content with the audience’s emotional needs by incorporating emotionally charged language.
Exaggerated (Ex) (exaggerate). This strategy uses exaggeration and strong rhetorical techniques to make the product appear more appealing.
Subliminal (Su) (subliminal). This strategy is a form of covert advertising that embeds information through subconscious cues. Although readers may not consciously recognize the advertising content, the implicit messages or psychological suggestions subtly influence their behavior.
We employ GPT-4o (gpt4o) to generate tool descriptions that exhibit specific advertising characteristics.
Algorithm for Descriptions Stealthiness Enhancement.
consists of two main components: advertising style transformation and genetic algorithm stealthiness enhancement. We first utilize GPT-4o (gpt4o) and advertising prompt to transform the original tool description into a style that aligns with the selected advertising strategy, while maintaining a certain level of stealthiness, after the initialization of the pool . Subsequently, a GA is applied to further enhance the stealthiness of the optimized description by iteratively refining candidate prompts. Specifically, in each iteration, we introduce the Mutate operation using the stealthiness-oriented prompt designed to improve stealthiness and perform the Crossover operation, combining elements from pairs of prompts to promote mutation diversity and explore a broader solution space. The resulting candidate descriptions are accumulated in a pool , from which the GPT-4o selects the top-k descriptions that appear least suspicious. These descriptions are retained for the next iteration, thereby guiding the evolutionary process toward higher stealthiness. After iterations, GPT-4o is used to select the most stealthy description from the final pool. The prompt design for each advertising strategy and the further details are shown in the Appendix.
5 Experimental Setup